Find the degree of the extension Q(V3+2V2) over Q.
Answer #1

Solution:

Your solution is here by preserved:

We started out by constructing polynomial f(x) = Q(2] such that 3+282 is a root of fa) x= 53+252 3+252 * x²=34252 (2²3)² (252

Any questions please comment.

Thank you